OpenStudy (anonymous):

f(x)=x^2 and g(x)=x+6 find g(f(0)) will fan and give medal

OpenStudy (zale101):

f(x)= x^2 g(x)= x+6 g(0) means the value of g(x) when x = 0

OpenStudy (anonymous):

than whats g(f)

OpenStudy (zale101):

g(0)= 0+6 so g(x)=6 plug that x=6 into f(x) and solve f(g(0)) = f(6 ) = 6^2

OpenStudy (anonymous):

that makes the answer 36

OpenStudy (zale101):

yes